OpenStudy (anonymous):

4. What are the points of discontinuity? Are they all removable? Please show your work.

OpenStudy (anonymous):

factor the denominator, then set it equal to zero and solve those are the discontinuities

OpenStudy (anonymous):

if, once you factor, you can also cancel, that will remove the discontinuity

OpenStudy (anonymous):

y = (x-5)/ x^2-6x+5

OpenStudy (anonymous):

can you help me step by step please

OpenStudy (anonymous):

factor the denominator

OpenStudy (anonymous):

\[y=\frac{x-5}{(x-5)(x-1)}\]

OpenStudy (anonymous):

the denominatior for my quation is x^2-6x+5 ight

OpenStudy (anonymous):

then if \[(x-5)(x-1)=0\] you have \[x=5,x=1\] and those are your discontinuities because you cannot divide by zero

OpenStudy (anonymous):

yes the bottom is the denominator

OpenStudy (anonymous):

okay can we got slow step by step like u ask me and i answer setc back and forth likethat please im having a hard time understanding and i want to learn

OpenStudy (anonymous):

how to i facto out that denominator is confusing to me im sorry :(

OpenStudy (anonymous):

ok do you know how to factor \[x^2-6x+5\]? because that is what you have to do next

OpenStudy (anonymous):

you do not "factor it out" you just factor it in to a product, as if you were solving the equation \[x^2-6x+5=0\]

OpenStudy (anonymous):

okay im not sure how to do that

OpenStudy (anonymous):

think of two numbers that, when multiplied together give \(5\) and when added together give \(-6\)

OpenStudy (anonymous):

2.5 x 2.5 = 5 right

OpenStudy (anonymous):

oh no think of two integers, not decimals

OpenStudy (anonymous):

also \[2.5\times 2.5\neq 5\]

OpenStudy (anonymous):

there are really only two pairs of integers that multiply to give \(5\)

OpenStudy (anonymous):

this is so hard lol im sorry for needing help

OpenStudy (anonymous):

no problem

OpenStudy (anonymous):

it is clear that \(1\times 5=5\) ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

and also \(-1\times -5=5\) right?

OpenStudy (anonymous):

may i ask why there are 3 equations like that ? and how each gets smaller

OpenStudy (anonymous):

yes that is correct

OpenStudy (anonymous):

but \(1+5=6\) and \(-1-5=-6\) so you have to pick \(-1,-5\) to factor

OpenStudy (anonymous):

okay so factor -1 and -5?

OpenStudy (anonymous):

why is there the 3 copies on the equations im confused by that

OpenStudy (anonymous):

factor as \[x^2-6x+5=(x-1)(x-5)\]

OpenStudy (anonymous):

if you multiply out on the right, you get the quadratic on the left

OpenStudy (anonymous):

there are not three copies, not sure what you mean by that

OpenStudy (anonymous):

ill screen shot for u

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

well idk how to on here so. is it just supposed to be x^2-6x+5 = (x-1)(x-5)

OpenStudy (anonymous):

yes those are equal

OpenStudy (anonymous):

okay good

OpenStudy (anonymous):

so \[\frac{x-5}{x^2-6x+5}=\frac{x-5}{(x-5)(x-1)}\]

OpenStudy (anonymous):

because the bottoms are equal

OpenStudy (anonymous):

now we are almost done

OpenStudy (anonymous):

you are looking for the "discontinuities" right?

OpenStudy (anonymous):

okay i understand

OpenStudy (anonymous):

yes the points of discontinuties and if they are removable

OpenStudy (anonymous):

you cannot divide by zero set \[(x-5)(x-1)=0\] and solve for \(x\)

OpenStudy (anonymous):

do you know how to do that?

OpenStudy (anonymous):

would it b 5-5 amd 1-1?

OpenStudy (anonymous):

it would be \(x=5\) or \(x=1\) so yeah, sort of

OpenStudy (anonymous):

\[x-5=0\\ x=5\] or \[x-1=0\\ x=1\]

OpenStudy (anonymous):

those are your discontinuities, \(5,1\)

OpenStudy (anonymous):

okay so 5,1 are my discontinuties

OpenStudy (anonymous):

yup now for the removable part

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

we have \[y=\frac{x-5}{(x-5)(x-1)}\] so you can cancel right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

what do you get when you cancel?

OpenStudy (anonymous):

0?

OpenStudy (anonymous):

hmm no

OpenStudy (anonymous):

can you explain it to me please

OpenStudy (anonymous):

\[y=\frac{x-5}{(x-5)(x-1)}=\frac{\cancel{x-5}}{\cancel{(x-5)}(x-1)}=\frac{1}{x-1}\]

OpenStudy (anonymous):

so x-1?

OpenStudy (anonymous):

or 1/x-1

OpenStudy (anonymous):

\[\frac{1}{x-1}\]

OpenStudy (anonymous):

okay great now what

OpenStudy (anonymous):

we "removed" the \(x-5\) right?

OpenStudy (anonymous):

correct

OpenStudy (anonymous):

so \(5\) is the "removable discontinuity"

OpenStudy (anonymous):

and 1 is not

OpenStudy (anonymous):

okay that makes sense, does the complete the answer?

OpenStudy (anonymous):

finally, yes

OpenStudy (anonymous):

okay, i appreciate the help so much. I have a few more ill be posting in a bit. hopefully your on. i love learning and just not getting an answer

OpenStudy (anonymous):

no problem